OpenStudy (anonymous):

Is anyone taking Algebra 2 on FLVS?

OpenStudy (anonymous):

What is FLVS??

OpenStudy (anonymous):

Florida Virtual School

OpenStudy (anonymous):

I live in florida too!

OpenStudy (anonymous):

okay. Are you taking this course?

OpenStudy (anonymous):

No :( im taking a different math course on E2020

OpenStudy (anonymous):

What do you want help with in Algebra 2???

OpenStudy (anonymous):

1.Solve S = 4v2 for v. 2.Solve M = 2x + 3y for y 3.Solve t = the quantity p plus 3 r, divided by six for r. 4.Solve V = π r2h for h.

OpenStudy (anonymous):

1.)v=+/- sqrt(s)/2

OpenStudy (anonymous):

2.) M=2x+3y Your goal should be to isolate y on one side M - 2x = 2x + 3y -2x Subtract 2x from both sides of the equation M - 2x = 3y (M-2x)/3=3y/3 Divide both sides of the equation by 3 (M-2x)/3=y y=M/3-2x/3 or y=(M-2x)/3

OpenStudy (anonymous):

3.)t=(P+3r)/6 6t=P+3r 6t-P=3r (6t-P)/3=r

OpenStudy (anonymous):

Solve S = 4v2 for v. divide both sides by 4 S/4 = v^2 now take square root of both sides +- sqrt(S/4) = v

OpenStudy (anonymous):

4.)h = V / (π r2)

OpenStudy (anonymous):

Thank You! :)

OpenStudy (anonymous):

thanks to who?

OpenStudy (anonymous):

You & the other person